# Finite-size spectrum of the staggered six-vertex model with antidiagonal boundary conditions

### v1

@@ Abstract
The finite-size spectrum of the critical staggered six-vertex model with antidiagonal boundary conditions is studied. Similar to the case of periodic boundary conditions, we identify three different phases. In two of those, the underlying conformal field theory can be identified to be related to the twisted $U(1)$ Kac-Moody algebra. In contrast, the finite size scaling in the third regime, whose critical behaviour with the quasi-periodic BCs is described by the $SL(2,\mathbb{R})_{k}/U(1)$ black hole CFT possessing a non-compact degree of freedom, is more subtle. Here with antidiagonal BCs imposed, the corrections to the scaling of the ground state grow logarithmically with the system size, while the energy gaps appear to close logarithmically. Moreover, we obtain an explicit formula for the Q-operator which is useful for numerical implementation.
@@ I Introduction and main results
In phase I, i.e. for staggering parameter $\alpha$ less than $\min(\gamma,\pi-\gamma)$ (or $\alpha>\max(\gamma,\pi-\gamma)$ related by duality, see Eq. (22) below), the critical behaviour is that of the homogeneous model, i.e. described by a free boson with compactification radius depending on $\gamma$. In the phases around the ’self-dual’ line $\alpha=\pi/2$ the field content of the low energy effective theory depends on the anisotropy: for $\gamma>\pi/2$ (phase II) the scaling limit of the model is described by a conformal field theory (CFT) consisting of a free compact boson and two Majorana fermions Ikhlef et al. (2010); Kotousov and Lukyanov (2021). Finally, in phase III realized for $\gamma<\pi/2$ the CFT is related to the $SL(2,\mathbb{R})_{k}/U(1)$ black hole coset model with a continuous component to the conformal spectrum Ikhlef et al. (2012); Candu and Ikhlef (2013); Frahm and Seel (2014); Bazhanov et al. (2021a).
@@
The appearance of such CFTs in the scaling limit of a lattice model with finitely many degrees of freedom per site has motivated further studies of this model with different boundary conditions (BCs). In a series of works integrable open boundary conditions leaving the lattice model invariant under the quantum group $U_{q}(\mathfrak{sl}(2))$ have been studied. Interestingly, two such boundary conditions can be constructed for the lattice model, both influencing the physical properties of the model in phase III: one of these BCs leads to a purely discrete set of conformal weights leading to a compact boundary CFT describing continuum limit of the model Robertson et al. (2019). With the second choice for the BCs the symmetry of the lattice model’s ground state is spontaneously broken. The finite size spectrum, however, contains both discrete and continuous parts allowing for the decomposition into irreps of the $\mathcal{W}_{\infty}$ algebra, the extended conformal symmetry of the model Robertson et al. (2021); Frahm and Gehrmann (2022); Frahm et al. (2023). While there remain some open questions regarding the relation between this scaling limit and possible D-brane constructions for the black hole CFT both numerical and analytical studies indicate that the two BCs realize an RG flow from an (unstable) non-compact boundary CFT to a (stable) compact one Robertson et al. (2021); Frahm and Gehrmann (2023).
@@
Here we continue these studies of the influence of BCs on the critical properties of the staggered six-vertex model by considering anti-diagonal BCs which break the continuous symmetry of the staggered six-vertex model into discrete ones. The homogeneous model with such BCs has been solved using Bethe ansatz methods Yung and Batchelor (1995); Batchelor et al. (1995) and the conformal spectrum in the different symmetry sectors has been found to correspond to a twisted $U(1)$ Kac-Moody algebra and does not depend on the anisotropy $\gamma$ von Gehlen et al. (1988); Alcaraz et al. (1988); Niekamp et al. (2009). Our paper is organized as follows: in the following section, we define the model and construct the commuting operators characterizing the spectrum and its symmetries. In Section III the finite-size spectrum of the model in the different phases is studied. To identify the root configurations parameterizing the low energy Bethe states we have generalized the construction of the Baxter Q-operator developed for the homogeneous model Batchelor et al. (1995) to the inhomogeneous case and present an explicit formula for its matrix elements in an appendix. Since the latter does not involve any matrix operations it is particularly suitable for an implementation on a computer. Based on these root configurations we construct RG trajectories to extract the scaling dimensions from the finite size data for the low-lying eigenenergies. We find that the low energy modes in phases I and II can be described in terms of twisted $U(1)$ Kac-Moody algebras, similar as in the homogeneous model von Gehlen et al. (1988); Alcaraz et al. (1988). In phase III, however, the corrections to scaling along the RG trajectories grow logarithmically with the system size. Such a behaviour has been observed before for particular states in the periodic model where they have been argued to leave the low energy spectrum and therefore not to be relevant for the scaling limit. Here all low energy states that we have identified for small system sizes show this behaviour. For the gap between the ground state and the first excitation, however, the corrections to scaling appear to close in the scaling limit. Larger system sizes need to be considered though to make a quantitative statement on the finite-size scaling. Therefore the characterization of the scaling limit in this phase is left to a future research project.
@@ II Definition of the model
The transfer matrix $\mathbbm{t}(u)$ acts on circular lattice of $2L$ sites i.e. the Hilbert space $\mathscr{H}$ is given by $\mathscr{H}=V_{1}\otimes\dots\otimes V_{2L}$. The influence of the twist matrix $\sigma^{x}_{0}$ is encoded in the boundary conditions: given an operator $B_{j}$ acting non-trivially on the $j^{\text{th}}$ site we identify
@@
This equation for $Q(u)$ can be solved by taking into account the analytic properties of the transfer matrix eigenvalues $t(u)$. As the transfer matrix is a Laurent polynomial in $\mbox{e}^{u}$, the asymptotic behaviour of its eigenvalues regarding the spectral parameter $u$ is given by
@@
Using (12) one can deduce from (11) that $Q(u)$ has to be of the the form:
@@
$-L\cot(\alpha-\gamma)+2L\cot(\gamma)+L\cot(\alpha+\gamma)$
It should be pointed out that the Hamiltonian is not self-adjoint for generic values of the anisotropy and the staggering. Moreover, the $U(1)$ charge $S^{z}=\sum_{j}\sigma_{j}^{z}$, which commutes with both the transfer matrix and the Hamiltonian for periodic boundary conditions, is broken to two discrete $\mathbb{Z}_{2}$ symmetries in the antidiagonal case considered here, see Ref. Alcaraz et al. (1988). They can be expressed as products over Pauli matrices
@@
By negation, we can neutralise the influence of the duality transformation $\mathcal{D}$ on the Hamiltonian. The resulting Hamiltonian (and its spectrum) is then invariant. In contrast, the eigenvalues of transfer matrix and the Q-operator will be modified since the transformation on the anisotropy $\gamma$ cannot be absorbed by the periodicity of the hyperbolic functions in (13)–(15). This leads to the fact that there exist two interchangeable sets of Bethe roots describing the spectrum of $\mathbbm{H}$; see for comparison the Figure 4. For technical reasons, the Bethe root patterns of the duality transformed roots are sometimes slightly more convenient for further study. Therefore, we always use the ‘dual’ root configurations in the following.
@@
Figure 4: The different Bethe root configurations for an excited state related by the transformation (23) for $162L=$ 16, $\gamma=2\pi/5$, $\alpha=\pi/2$: in both cases the roots tend to align on lines with constant imaginary part. The deviations from these lines, however, are smaller for the dual root configuration (right panel) which makes this parameterization more convenient for the finite size analysis.
@@
is called the quasi-momentum operator. It plays an important role in identifying the scaling limit of the staggered six-vertex model for the boundary conditions studied in previous works Ikhlef et al. (2008, 2012); Bazhanov et al. (2021a, b); Frahm and Gehrmann (2022); Robertson et al. (2020); Frahm and Gehrmann (2023); Frahm et al. (2023).
@@ III Finite-size studies
In the following sections, we investigate the low energy spectrum of the lattice model for increasing system sizes to obtain information about the effective field theory arising in its scaling limit. For the case at hand, an integrable lattice model, this task is facilitated by the description of individual states in terms of their corresponding Bethe root configurations: consider the model with $L_{\rm in}$ sites and a state |$\ket{\Psi_{L_{\rm in}}}$ with Bethe roots $\{v_{k}\}^{L_{\rm in}}$. If the pattern of the Bethe roots $\{v_{k}\}^{L_{\rm in}+2}\,,\{v_{k}\}^{L_{\rm in}+4},\dots$ of states |$\ket{\Psi_{L_{\rm in}+2}}\,,\ket{\Psi_{L_{\rm in}+4}},\dots$ of larger systems sizes $L_{\rm in}+2\,,L_{\rm in}+4,\dots$, is qualitatively the same as the one of $\{v_{k}\}^{L_{\rm in}}$, these states can be grouped together in one so-called RG-trajectory {|$\{\ket{\Psi_{L}}\}$. We can construct the RG-trajectory {|$\{\ket{\Psi_{L}}\}$ up to $L\sim 2000$ by merely solving the Bethe ansatz equation without relying on a direct diagonalization of the Hamiltonian. The latter is an impossible task for as the size of the Hilbert space grows exponentially with the system size $L$.
As a starting point for the construction of these RG trajectories, one needs the Bethe root configurations for small system sizes. For small $L$ these can be obtained by a simultaneous diagonalization of the Hamiltonian and the Q-operator as the Bethe roots are the zeros of $Q(u)$, see (13). In this way, we can extract the Bethe root configurations of the first few hundred low-energy states. These are then taken as the seed for the construction of the corresponding RG trajectories via the procedure described above.
@@ III.1 Phase I
In terms of the real parts $x_{k},y_{j}$, the logarithmic Bethe equations for the ground state reads
@@
The majority of the Bethe root of a low energy excitation above the ground state is still given by (30) while some roots are located in the complex plane either as strings or single complex roots. Our numerical work shows that the effective scaling dimensions are given by
@@
which is twice the partition function of a CFT built from the (unique) irreducible representation of a twisted $U(1)$ Kac-Moody algebra, one for each sector of the $\mathbb{Z}_{2}$-symmetry $\mathbb{G}$ (20). The central charge is one ($c=1$) and the zero mode of the Virasoro algebra takes the form von Gehlen et al. (1988)
@@ III.2 Phase II
Note that this is the square of the corresponding function (37) in phase I. As a result we propose that the CFT describing the scaling limit of the staggered vertex model in phase II has a central charge $c=2$ and is built from two copies of the twisted $U(1)$ Kac-Moody algebra. Here, the zero mode of the Virasoro algebra takes the form
@@
For periodic boundary conditions, the scaling limit of the staggered model in this phase was shown to be related to one compact boson and two Majorana fermions Ikhlef et al. (2010); Kotousov and Lukyanov (2021). Ad hoc, the present findings for antidiagonal BC do not contradict the results of the staggered model with periodic ones. This is due to the fact that the modules of the Majoranas will start to interfere with the ones of the boson once the antiperiodicity is imposed. This can lead to a restructuring of all the modules, which will then be equivalent to the case at hand.
@@ III.3 Phase III
Bulk quantities, such as the energy density of a spin chain in the thermodynamic limit, do not depend on the specific choice boundary conditions imposed. We have already seen this phenomenon in phases I and II. Hence, we expect that in phase III, the energy density will be identical to the known one of the periodic model Frahm and Seel (2014):
@@
Figure 6: Lowest scaling dimensions (29) in phase III: red circles are derived from the lowest real eigenvalues of (19) for $L=10$, i.e. $20$ lattice sites. Black crosses from the real part of complex eigenvalues.
Moreover, the study of small systems does not yield reliable support for the construction of RG trajectories based our Bethe hypothesis (48) as the root patterns may change drastically when $L$ is varied, see Figure 7.
@@
for Bethe root configuration displayed in the right panel of Figure 4.
@@
The energy gaps appear to close as $\sim{1}/{L\log(L)}$. Further, the closing of the energy gap seems to be linearly related to the quasi-momentum operator (26). This might signal the remains of the continuous component of the finite-size spectrum observed in phase III for quasi-periodic BC, although the scaling of the ground state is very different. An investigation of this peculiar behaviour in more detail requires an extensive systematic numerical study. As discussed above, such a study faces the difficulty that small $L$ solutions to the Bethe equations often can not be used to seed the corresponding RG-trajectories due to a significant number of level crossings with the system size. Moreover, the construction of the RG-trajectories of these states turns out to be very time-consuming since the root-finding algorithms appear to be extremely sensitive to the initial values. Therefore, we leave the characterization of the scaling limit in this phase as a future research project on its own. Our numerical data for the Bethe root configurations of the ground state and first excitation used in Figures 8, 9 and 10 are available online Frahm and Gehrmann (2024).

### v2

@@ Abstract
The finite-size spectrum of the critical staggered six-vertex model with antidiagonal boundary conditions is studied. Similar to the case of periodic boundary conditions, we identify three different phases. In two of those, the underlying conformal field theory can be identified to be related to the twisted $U(1)$ Kac-Moody algebra. In contrast, the finite size scaling in the third regime, whose critical behaviour with the (quasi-)periodic BCs is related to the 2d black hole CFTs possessing a non-compact degree of freedom, is more subtle. Here with antidiagonal BCs imposed, the corrections to the scaling of the ground state grow logarithmically with the system size, while the energy gaps appear to close logarithmically. Moreover, we obtain an explicit formula for the Q-operator which is useful for numerical implementation.
@@ I Introduction and main results
In phase I, i.e. for staggering parameter $\alpha$ less than $\min(\gamma,\pi-\gamma)$ (or $\alpha>\max(\gamma,\pi-\gamma)$ related by duality, see Eq. (22) below), the critical behaviour is that of the homogeneous model, i.e. described by a free boson with compactification radius depending on $\gamma$. In the phases around the ’self-dual’ line $\alpha=\pi/2$ the field content of the low energy effective theory depends on the anisotropy: for $\gamma>\pi/2$ (phase II) the scaling limit of the model is described by a conformal field theory (CFT) consisting of a free compact boson and two Majorana fermions Ikhlef et al. (2010); Kotousov and Lukyanov (2021). Finally, in phase III realized for $\gamma<\pi/2$ the critical behaviour is related to the 2d black hole CFTs which possess a continuous component in the conformal spectrum Ikhlef et al. (2012); Candu and Ikhlef (2013); Frahm and Seel (2014); Bazhanov et al. (2021a).
@@
The appearance of such CFTs in the scaling limit of a lattice model with finitely many degrees of freedom per site has motivated further studies of this model with different boundary conditions (BCs). In a series of works integrable open boundary conditions leaving the lattice model invariant under the quantum group $U_{q}(\mathfrak{sl}(2))$ have been studied. Interestingly, two such boundary conditions can be constructed for the lattice model, both influencing the physical properties of the model in phase III: one of these BCs leads to a purely discrete set of conformal weights leading to a compact boundary CFT describing the continuum limit of the model Robertson et al. (2019). With the second choice for the BCs the symmetry of the lattice model’s ground state is spontaneously broken. The finite size spectrum, however, contains both discrete and continuous parts allowing for the decomposition into irreps of the $\mathcal{W}_{\infty}$ algebra, the extended conformal symmetry of the model Robertson et al. (2021); Frahm and Gehrmann (2022); Frahm et al. (2023). While there remain some open questions regarding the relation between this scaling limit and possible D-brane constructions for the black hole CFT both numerical and analytical studies indicate that the two BCs realize an RG flow from an (unstable) non-compact boundary CFT to a (stable) compact one Robertson et al. (2021); Frahm and Gehrmann (2023).
@@
Here we continue these studies of the influence of BCs on the critical properties of the staggered six-vertex model by considering anti-diagonal BCs which break the $U(1)$ symmetry of the staggered six-vertex model into discrete ones. The homogeneous model with such BCs has been solved using Bethe ansatz methods Yung and Batchelor (1995); Batchelor et al. (1995) and the conformal spectrum in the different symmetry sectors has been found to correspond to a twisted $U(1)$ Kac-Moody algebra and does not depend on the anisotropy $\gamma$ von Gehlen et al. (1988); Alcaraz et al. (1988); Niekamp et al. (2009). Our paper is organized as follows: in the following section, we define the model and construct the commuting operators characterizing the spectrum and its symmetries. In Section III the finite-size spectrum of the model in the different phases is studied. To identify the root configurations parameterizing the low energy Bethe states we have generalized the construction of the Baxter Q-operator developed for the homogeneous model Batchelor et al. (1995) to the inhomogeneous case and present an explicit formula for its matrix elements in an appendix. Since the latter does not involve any matrix operations it is particularly suitable for an implementation on a computer. Based on these root configurations we construct RG trajectories to extract the scaling dimensions from the finite size data for the low-lying eigenenergies. We find that the low energy modes in phases I and II can be described in terms of twisted $U(1)$ Kac-Moody algebras, similar as in the homogeneous model von Gehlen et al. (1988); Alcaraz et al. (1988). In phase III, however, the corrections to scaling along the RG trajectories grow logarithmically with the system size. Such a behaviour has been observed before for particular states in the periodic model where they have been argued to leave the low energy spectrum and therefore not to be relevant for the scaling limit. Here all low energy states that we have identified for small system sizes show this behaviour. For the gap between the ground state and the first excitation, however, the corrections to scaling appear to close in the scaling limit. Larger system sizes need to be considered though to make a quantitative statement on the finite-size scaling. Therefore the characterization of the scaling limit in this phase is left to a future research project.
@@ II Definition of the model
The transfer matrix $\mathbbm{t}(u)$ acts on a circular lattice of $2L$ sites i.e. the Hilbert space $\mathscr{H}$ is given by $\mathscr{H}=V_{1}\otimes\dots\otimes V_{2L}$. The influence of the twist matrix $\sigma^{x}_{0}$ is encoded in the boundary conditions: given an operator $B_{j}$ acting non-trivially on the $j^{\text{th}}$ site we identify
@@
This equation for $Q(u)$ can be solved by taking into account the analytic properties of the transfer matrix eigenvalues $t(u)$. The transfer matrix is a Laurent polynomial in $\mbox{e}^{u}$. The asymptotic behaviour of its eigenvalues regarding the spectral parameter $u$ is given by
@@
Using (12) one can deduce from (11) that $Q(u)$ has to be of the form:
@@
$-L\cot(\alpha-\gamma)+2L\cot(\gamma)+L\cot(\alpha+\gamma)\,.$
It should be pointed out that the Hamiltonian is not self-adjoint for generic values of the anisotropy and the staggering. Moreover, the global $O(2)$ symmetry of the model is broken to two discrete $\mathbb{Z}_{2}$ symmetries in the antidiagonal case considered here, see Ref. Alcaraz et al. (1988). They can be expressed as products over Pauli matrices
@@
By negation, we can neutralise the influence of the duality transformation $\mathcal{D}$ on the Hamiltonian. The resulting Hamiltonian (and its spectrum) is then invariant. In contrast, the eigenvalues of the transfer matrix and the Q-operator are periodic functions in $\gamma$ but with period $4\pi$ (e.g. see (13)). Hence, they are modified under the duality transformation. This leads to the fact that there exist two interchangeable sets of Bethe roots describing the spectrum of $\mathbbm{H}$; see for comparison the Figure 4. For technical reasons, the Bethe root patterns of the duality transformed roots are sometimes slightly more convenient for further study. Therefore, we always use the ‘dual’ root configurations in the following.
@@
Figure 4: The different Bethe root configurations for an excited state related by the transformation (23) for $162L=$ 16, $\gamma=2\pi/5$, $\alpha=\pi/2$: in both cases the roots tend to align on lines with constant imaginary part. The deviations from these lines, however, are smaller for the dual root configuration (right panel) which makes this parameterization more convenient for the finite size analysis (see Frahm and Gehrmann (2024) for the numerical data used in this figure).
@@
is called the quasi-momentum operator. It plays an important role in identifying the scaling limit of the staggered six-vertex model for the boundary conditions studied in previous works Ikhlef et al. (2012); Bazhanov et al. (2021a, b); Frahm and Gehrmann (2022); Robertson et al. (2020); Frahm and Gehrmann (2023); Frahm et al. (2023).
@@ III Finite-size studies
In the following sections, we investigate the low energy spectrum of the lattice model for increasing system sizes to obtain information about the effective field theory arising in its scaling limit. For the case at hand, an integrable lattice model, this task is facilitated by the description of individual states in terms of their corresponding Bethe root configurations $\{v_{k}\}^{2L}$: starting from a particular eigenstate |$\ket{\Psi_{L_{\rm in}}}$ of the model with $L_{\rm in}$ sites we construct so-called RG-trajectories {|$\{\ket{\Psi}_{L}\}$ by grouping states |$\ket{\Psi_{L}}$ with $L=L_{\rm in},L_{\rm in}+2,L_{\rm in}+4,\dots$ with Bethe roots forming qualitatively similar patterns, see Figure 5 for an illustration.
[…]
Figure 5: Construction of an RG-trajectory: shown are the dual Bethe root configurations of particular states for $\gamma=\frac{\pi}{6}$, $\alpha=\frac{\pi}{2}$ for $162L=$ 16 (left panel) and $20$ (right panel): the states are identified as members of the same RG-trajectory since the root patterns are characterized by their main features, i.e. two purely imaginary roots at $\frac{{\rm i}\pi}{6}$ and $-\frac{2{\rm i}\pi}{3}$ apart from the bulk of roots with imaginary parts near $0,\pm\frac{\pi}{2},-\pi$.
This allows to construct RG-trajectories {|$\{\ket{\Psi_{L}}\}$ for system sizes up to $L\sim 2000$ by merely solving the Bethe ansatz equation without relying on a direct diagonalization of the Hamiltonian. The latter is an impossible task as the size of the Hilbert space grows exponentially with the system size $L$.
The initial data for this construction, i.e. the Bethe root configurations for some small system size $L_{\rm in}$, are obtained by a simultaneous diagonalization of the Hamiltonian and the Q-operator. The root configurations for the few hundred lowest energy states are obtained by finding the zeros of $Q(u)$, see (13). These are then taken as the seed for the construction of the corresponding RG trajectories via the procedure described above.
@@ III.1 Phase I
In terms of the real parts $x_{k},y_{j}$, the logarithmic Bethe equations for the ground state read
@@
The majority of the Bethe roots of a low energy excitation above the ground state is still given by (30) while some roots are located in the complex plane either as strings or single complex roots. Our numerical work shows that the effective scaling dimensions are given by
@@
We have checked this up to the indicated order numerically. The degeneracy formula (37) corresponds to twice the partition function of a CFT built from the (unique) irreducible representation of a twisted $U(1)$ Kac-Moody algebra, one for each sector of the $\mathbb{Z}_{2}$-symmetry $\mathbb{G}$ (20). The central charge is one ($c=1$) and the zero mode of the Virasoro algebra takes the form von Gehlen et al. (1988)
@@ III.2 Phase II
We have checked this up to the indicated order numerically. Note that this is the square of the corresponding function (37) in phase I. As a result we propose that the CFT describing the scaling limit of the staggered vertex model in phase II has a central charge $c=2$ and is built from two copies of the twisted $U(1)$ Kac-Moody algebra. Here, the zero mode of the Virasoro algebra takes the form
@@
For periodic boundary conditions, the scaling limit of the staggered model in this phase was shown to be related to one compact boson and two Majorana fermions Ikhlef et al. (2010); Kotousov and Lukyanov (2021). Note that the present findings for antidiagonal BC do not contradict the results of the staggered model with periodic ones. This is due to the fact that the modules of the Majoranas will start to interfere with the ones of the boson once the antiperiodicity is imposed. This can lead to a restructuring of all the modules, which will then be equivalent to the case at hand.
@@ III.3 Phase III
Bulk quantities, such as the energy density of a spin chain in the thermodynamic limit, do not depend on the specific choice of boundary conditions imposed. We have already seen this phenomenon in phases I and II. Hence, we expect that in phase III, the energy density will be identical to the known one of the periodic model Frahm and Seel (2014):
@@
Figure 7: Lowest scaling dimensions (29) in phase III: red circles are derived from the lowest real eigenvalues of (19) for $\alpha=\frac{\pi}{2}$ and $L=10$, i.e. $20$ lattice sites. Black crosses from the real part of complex eigenvalues.
Moreover, the study of small systems does not yield reliable support for the construction of RG trajectories based on our Bethe hypothesis (48) as the root patterns may change drastically when $L$ is varied, see Figure 8.
@@
for the Bethe root configuration displayed in the right panel of Figure 4.
@@
The energy gap appears to close as $\sim{1}/{L\log(L)}$. Further, the closing of the energy gap seems to be linearly related to the quasi-momentum operator (26). This might signal the remains of the continuous component of the finite-size spectrum observed in phase III for quasi-periodic BC, although the scaling of the ground state is very different.
In summary, taking the RG trajectories seeded by the ground state and first excitation for $L=8$ as a reference the corrections to scaling diverge while the energy gap to the first excitation closes logarithmically:
$L(E_{0}-L\epsilon_{\infty})$ $\sim a_{1}+a_{2}\log^{2}(L)\,,$ (52)
$L(E_{1}-E_{0})$ $\sim\tilde{a}_{1}+\frac{\tilde{a}_{2}}{\log(L)}\,,$
where $a_{i},\,\tilde{a}_{i}\in\mathbb{C}$. Speculating about the origin of the logarithmic growth of the subleading corrections to $E_{0}$ one may consider two possible scenarios: (a) the corresponding RG trajectory does not connect to the true ground state (or even belong to the low energy sector) with bulk energy (47) in the scaling limit. (b) The RG trajectory does describe the ground state of the model but the antidiagonal boundary conditions lead to a marginal relevant perturbation of the CFT describing the scaling limit for quasi-periodic boundary conditions (with diagonal twist). As for (a) we have not observed any evidence for the existence of another set of RG trajectories starting from the small $L$ low energy spectrum with decreasing corrections to scaling as $L\to\infty$. Hence, to rule out the existence of such a set the finite size analysis of the spectrum has to be extended significantly. As discussed above, such an analysis has to address the problem of a large number of level crossings (with the system size) interfering with the construction of the corresponding RG trajectories. In the case (b) the finite size scaling (52) describes the RG flow in the vicinity of the (unstable) fixed point under the perturbation by the antiperiodic boundary conditions. Since the integrable model does not contain any further parameter, however, it will not be possible to identify the corresponding stable fixed point within this setting.
In both cases the characterization of the scaling limit in phase III requires substantial additional effort. Therefore, we leave this problem for a future research project on its own. Our numerical data for the Bethe root configurations of the ground state and first excitation used in Figures 9, 10 and 11 are available online Frahm and Gehrmann (2024).